OpenStudy (booksforlife):

I'm not sure how to work this problem out. Solve open parentheses square root of 6 close parentheses to the 8 x power = 216x−3 Please help.

OpenStudy (mathmale):

\[(\sqrt{6})^8=216x - 3?\]

OpenStudy (booksforlife):

Yes

OpenStudy (mathmale):

Pls verify correctness of my previous entry.

OpenStudy (mathmale):

Now, simplify the radical:\[(\sqrt{6})^8=?\]

OpenStudy (mathmale):

Note that this is th

OpenStudy (mathmale):

this is the same as \[((\sqrt{6})^2)^4=?\]

OpenStudy (booksforlife):

The equation you did the first time looks like the one I have. I'm just not sure on how to solve it.

OpenStudy (mathmale):

What's the square of the square root of 6?

OpenStudy (booksforlife):

2.4?

OpenStudy (booksforlife):

Oh @mathmale I noticed in your equation you drew you didn't put 8x in the power. You just put 8.

OpenStudy (mathmale):

That's an indication that you need to use parentheses to show that the '8' and the 'x' go together. Please type in the original problem again, using parentheses so that there is no possibility of misinterpretation.

OpenStudy (booksforlife):

Okay, sorry I'm new at this. 6^(8x)=216^(x-3)

OpenStudy (booksforlife):

@mathmale is this better?

OpenStudy (mathmale):

That's a totally different problem, which emphasizes why you need to use parentheses whenever there's any possibility of doubt as to what you mean.\[6^{8x}=216^{(x-3)}\]

OpenStudy (mathmale):

Is 216 an integer power of 6? If so, write 6^x=216. Show your work, please.

OpenStudy (booksforlife):

I'm not sure about what your asking. :(

OpenStudy (mathmale):

6 raised to what power equals 216?

OpenStudy (booksforlife):

All I know is that 6 and 216 are the base and I need them to be the same. Once they are the same I take those away and just focus on the powers, but the radical sign is throwing me off. I am so confused! Ugh! I forgot the radical sign in my equation! I am so so so sorry! There should be a radical with the 6.

OpenStudy (mathmale):

Please, make the effort to answer my question about 6 and 216. First I asked you whether 216 is an integer power of 6. It is. Next I asked you "6 raised to what power equals 216?" Try raising 6 to the 1st, 2nd, 3rd, 4th and 5th powers. That might help establish the necessary pattern in your mind.

OpenStudy (booksforlife):

Okay would it be 6^3?

OpenStudy (mathmale):

Yes. Now re-write my most recent equation as is, except, replace "216" with "(6^3)".

OpenStudy (booksforlife):

So it would be 6^(8x)=6^(3)(x-3) with x-3 still in the power?

OpenStudy (mathmale):

Yes. Now you have the same base on both sides of your equation. What is it? Equate the power on the left side to the power on the right side.

OpenStudy (booksforlife):

I think now I just have to use the powers, but I need to distribute the 3. Is that correct?

OpenStudy (mathmale):

Yes. I'd rather that you plunge ahead and try things like this. You don't need my approval.

OpenStudy (mathmale):

Left side: expo is 8x. Right side: expo is ..... ?

OpenStudy (booksforlife):

Now we have 8x=3x-9?

OpenStudy (mathmale):

Yes. Again, don't wait for my approval. Go ahead / plunge ahead and find x; we will verify whether or not your result is correct or not.

OpenStudy (mathmale):

Your result: x = ??

OpenStudy (booksforlife):

-3?

OpenStudy (mathmale):

If 8x=3x-9, how would x come out equal to -3? Have I made any math errors? or have you?

OpenStudy (booksforlife):

I think I did. Let me try again.

OpenStudy (booksforlife):

Would I do this: 8x=3x-9 8x-3x = 3x-3x - 9 5x=-9 5x/5 = -9/5 x=-1.8?

OpenStudy (mathmale):

Here's where you'll need to check your own work. You began with 8x = 3x - 9. If you substitute x =-1.8, is the resulting equation true or not?

OpenStudy (booksforlife):

It is true!

OpenStudy (mathmale):

So, what is your conclusion?

OpenStudy (booksforlife):

x=-1.8 8x=3x-9 8(-1.8)=3(-1.8)-9 -14.4=-14.4

OpenStudy (mathmale):

Indeed. But I was askin you, "What does the fact that this check-up equation is true tell you about your solution to the original problem?"

OpenStudy (booksforlife):

I'm not sure what you mean by that. :(

OpenStudy (booksforlife):

I think I've got it now. Thank you!

OpenStudy (mathmale):

You've checked your result and found that your "check equation" was true. That means y our solution to the original problem is most likely correct. Please review what I said about using parentheses for enhanced clarity. Good work. See you again.